OpenStudy (anonymous):

use addition method. 3x+6y=0 4x-2y=30

OpenStudy (anonymous):

i got 7x-4y=30....now what?

OpenStudy (anonymous):

ur are correct whats the problem??? now???

OpenStudy (anonymous):

what is x and y ? (x,y) ?

OpenStudy (anonymous):

x is equal to 30 / 7 and y is equal to negatice 15/2

OpenStudy (anonymous):

2y=4x-30 3x+4x-30=0 7x-30=0 7x=30 x=30/7 4x30/7-2y=30 . . . 15/2

OpenStudy (anonymous):

negative 15/2 thats its

OpenStudy (anonymous):

click the best response please

OpenStudy (anonymous):

if u have done it with addition and u get 8x+4y=30 then upon subtracting the same equations u get x-8y=30 and now solve both these equation